OpenStudy (youngster):

Find x in this equation: 3+2cos x= 5 sin 2x

OpenStudy (youngster):

I got up to something like 1.5=cos x (5sin x -1)

OpenStudy (anonymous):

When you're done @satellite73 could you please help me with my DE problem?

OpenStudy (anonymous):

lets try \[5\sin(2x)-2\cos(x)-3=0\] \[10\sin(x)\cos(x)-2\cos(x)-3=0\] hmm that doesn't seem to help

OpenStudy (anonymous):

@MathSofiya i would love to but i really forgot the very few things i knew about ODE

OpenStudy (youngster):

sin (a+b) = sin a cos b - cos a sin b , right?

OpenStudy (youngster):

@satellite73 ?

OpenStudy (callisto):

It's a + sin (a+b) = sin a cos b + cos a sin b

OpenStudy (youngster):

I meant -

OpenStudy (youngster):

(subtraction) That is what I used in my work.

OpenStudy (callisto):

._.! Alright.. \[sin(a\pm b) = sina\ cosb \pm cosa \ sinb\]

OpenStudy (youngster):

But I still cannot figure out how to get x from\[1.5= \cos x (5 \sin x - 1)\]